# Simulated optical molasses cooling of trapped antihydrogen

### v1

@@ II Energy levels in magnetic field
In the ALPHA experiment, $\bar{\rm H}$s are laser cooled in a magnetic trap with a minimum $B\sim 1$ TBaker et al (ALPHA collaboration). This means the energy levels are strongly changed from their low (or zero) field character. Reference Rasmussen et al. (2017) discusses the energy levels. There are two 1S states that can be trapped called the 1Sc and 1Sd states. The 1Sd state has the positron and antiproton spin aligned; in the 1Sc state, they are antiparallel. To prevent losses from the photon emission step, the laser will excite the 2Pa state which has the positron spin and orbital angular momentum aligned to give total positron angular momentum $J=3/2,M=3/2$. Because of the different magnetic moments, the transition frequency for the 1Sc-2Pa transition is approximately 675 MHz higher than that for the 1Sd-2Pa transition at 1 T. The difference in frequencies does not change much with increasing $B$; for example, at $B=$0.5 T, the difference in frequencies is 660 MHz while the difference in frequencies is 680 MHz at 1.5 T.
@@ III Monte Carlo: no magnetic trap
As discussed in the introduction, laser cooling of trapped $\bar{\rm H}$ is complicated by the changing transition frequency due to the motion through the spatially varying $B$-field. The cooling is also complicated by the fact that the photons only travel in one direction and thus can only cool one component of the velocity. Cooling of all directions relies on the motion through the trap to mix the velocity components. In addition to the extra considerations from the constraints of the ALPHA trap, we found that optical molasses cooling of $\bar{\rm H}$ is complicated even when the transition frequency is constant and the photons can travel in any direction. This complication arises from the relatively large velocity kick that each photon absorption and emission imparts to the atom. Also, the recoil part of the excited state energy is sizable enough to have quantitative effects.
@@ III.1 Energy shift
A recoil term is usually left out due to smallness in determining the energy conservation of the transition. For $\bar{\rm H}$ cooling, this term is large enough that it could cause some changes in the results. Consider the energy before and after the photon absorption
@@ III.2 Ideal Optical Molasses
In this section, we describe Monte Carlo simulations of laser cooling of $\bar{\rm H}$ using an ideal optical molassesFoot (2005). Ideal optical molasses cooling assumes the photon has a wave vector, $\vec{k}$, in a random direction relative to the atom’s velocity, $\vec{v}$. The Doppler effect leads to an effective detuning of
@@
Our Monte Carlo simulation of the ideal optical molasses used the following algorithm: (1) for each of $N$ atoms, pick a random initial photon direction, $\vec{k}$, and for that atom compute the effective detuning, Eq. (3); (2) compute the probability that the photon was absorbed using $P=P_{0}/[1+(2\Delta/\Gamma)^{2}]$ with $P_{0}$ a small number of order 0.01; (3) compute a random number; if it is smaller than $P$, then a photon was absorbed; (4) if a photon is absorbed, compute a random emission direction and use Eq. (4) to update that atom’s velocity. If we wanted to simulate the effect of thermalization from elastic collisions between the atoms, we added a step: (5) randomly pick pairs of atoms and rotate their relative velocity to emulate a collision; repeated often enough step (5) leads to a Maxwell-Boltzmann velocity distribution at temperature $k_{B}T=(2/3)\langle KE\rangle$.
@@
Figure 1: For an ideal optical molasses, the steady state temperature using the Monte Carlo method versus the lab frame detuning in units of $\Gamma$. For (a), there is no collision between the atoms to achieve a Maxwell-Boltzmann velocity distribution. The different lines are for different $\bar{\rm H}$ masses: blue long dash ($1M$), green dash ($2M$), black dot ($4M$), orange dash-dot ($8M$), maroon dash-dot-dot ($16M$), purple dash-dot-dot-dot ($32M$), and red solid is the ideal case, Eq. (5). For (b), we compare the effect that elastic collisions leading to thermalization has on the steady state temperature: no collisions [blue long dash ($1M$), green dash ($2M$)], with collisions [orange dash-dot ($1M$), maroon dash-dot-dot ($2M$)].
The results of the simulations are shown in Fig. 1. Figure 1(a) shows the results of the simulations when we did not include collisions between the atoms, skipping step (5) of the algorithm. This figure shows the results when the photon wavelength is kept at 121.6 nm but we artificially change the mass of the $\bar{\rm H}$ from 1 to 32 times the actual mass by factors of 2. The blue long-dash curve is for the actual $\bar{\rm H}$ mass. The red solid curve is the ideal case, Eq. (5).
@@ IV Monte Carlo: antihydrogen trap
Some of the features specific for the optical molasses cooling include the path of the laser which is a straight line with a $2.3^{\circ}$ tilt relative to the trap axis. Thus, the $\vec{k}$ is a constant. The motion of the $\bar{\rm H}$ through the trap mixes the different velocity components possibly allowing for three dimensional cooling. Since this mixing is relatively slow, the detuning and intensity need to be chosen to limit the heating of the radial motion due the photon emission step. The beam has a waist of 3.48 mm and is pulsed with a duration of 10’s of nanoseconds and with repetition rate of 50 Hz. The laser cooling happens over a time scale measured in hours which is possible due to the cryogenic vacuum.
@@
One of the important features is that the detuning of the photon depends on the position in the trap as well as the Doppler shift. The reason for this is the $B$-field changes with position in the trap and the $1S-2P$ resonance frequency increases with increasing $B$-field. For negative detuning, the detuning becomes increasingly negative with increasing $B$-field. Figure 2 shows the magnitude of the $B$-field on the trap axis as a function of the position on the axis relative to the trap center. There are 5 mirror coils and 2 solenoidal coils that are used to shape the $B$-field on the axis. The red solid line is a flat field similar to that used in the 1S2S spectroscopy measurements in Ref. Ahmadi et al (ALPHA collaboration). The blue long dash line gives a more harmonic trap used in some of the simulations below. The green dashed and black dot curves are the slightly dipped and raised traps of Secs. IV.2.
@@
The laser line width, estimated as 60 MHz full-width half-maximum (FWHM), also modestly affects the results since the intrinsic linewidth is approximately 100 MHz. The laser line width is incorporated into the simulation by having each pulse at a slightly different detuning randomly chosen from a Gaussian distribution with a FWHM of 60 MHz. This leads to an effectively broader transition although the effect is not large. In all of the figures and results, the linewidth, $\Gamma$, is the natural radiative linewidth of the 2P state.
@@ IV.3 Cooling in a nearly harmonic trap
The harmonic trap, blue long dash curve in Fig. 2, has only a tiny region where the detuning has smallest magnitude. This suggests that the problems with the photon scattering rate being higher than the coordinate mixing rate will be less. The actual case is that the harmonic trap has much smaller mixing rate between the coordinates. We find that the temperature differences are much more extreme than the previous section. For example, at $\Delta_{0}=-3\Gamma$ and 2 nJ, $T_{z}\simeq 6$ mK and $T_{r}\simeq 24$ mK for the harmonic trap while the flat trap has $\simeq 6$ and $\simeq 9$ mK. Thus, nearly harmonic traps appear to be a poor choice for laser cooling of only the axial motion.
@@ V Simultaneous cooling 1Sc and 1Sd
Unfortunately, problems arise when laser cooling both populations and trying to achieve the same distributions for each. The 1Sc-2Pa transition frequency is 675 MHz larger at $B\sim 1$ T than the 1Sd-2Pa. Detuning to -250 MHz for the 1Sd-2Pa transition gives -925 MHz detuning for the 1Sc-2Pa transition. Thus, the cooling is extremely slow for the 1Sc states and the final distribution is at a substantially larger temperature with, essentially, uncooled $\bar{\rm H}$s.Baker et al (ALPHA collaboration)

### v2

@@ I Introduction
There are three main updates to Ref. Donnan et al. (2013). First, we recognize that some of the changes from optimal detuning are due to the small mass and large frequency of the cooling transition (Sec. III). Second, we use magnetic fields and laser parameters more representative of the ALPHA experiment (Sec. IV) leading to lower simulated temperatures. Lastly, we describe a possible method for simultaneously laser cooling the 1Sc and 1Sd trapped populations (Sec. V) leading to an important improvement to precision measurements.
@@ II Energy levels in magnetic field
In the ALPHA experiment, $\bar{\rm H}$s are laser cooled in a magnetic trap with a minimum $B\sim 1$ TBaker et al (ALPHA collaboration). This means the energy levels are strongly changed from their low (or zero) field character. Reference Rasmussen et al. (2017) discusses the energy levels. The energies in a magnetic field are labeled with a Roman letter at the end which, by convention, increases from low to high energy for 1S states while increasing from high to low energy for the 2P states (as ordered for small $B$). There are two 1S states that can be trapped called the 1Sc and 1Sd states. The 1Sd state has the positron and antiproton spin aligned; in the 1Sc state, they are antiparallel. To prevent losses from the photon emission step, the laser will excite the 2Pa state which has the positron spin and orbital angular momentum aligned to give total positron angular momentum $J=3/2,M=3/2$. Because of the different magnetic moments, the transition frequency for the 1Sc-2Pa transition is approximately 675 MHz higher than that for the 1Sd-2Pa transition at 1 T. The difference in frequencies does not change much with increasing $B$; for example, at $B=$0.5 T, the difference in frequencies is 660 MHz while the difference in frequencies is 680 MHz at 1.5 T.
The hyperfine splitting of the 2P states is several 10’s MHz compared to the order 10 GHz splitting of the states. Thus, the antiproton spin is nearly decoupled from the positron total angular momentum, $J$, at the $\sim 1$ T of the ALPHA experiment. The 1Sd-2Pa transition is exactly closed because all of the angular momenta are aligned. The 1Sc-2Pa transition is not exactly closed because the 1S state, with a more than 10×10$\times$10 × larger hyperfine splitting than 2P, has more mixing of the wrong direction antiproton spin. After the 1Sc-2Pa transition, the 2Pa state decays to the 1Sa state with a branching ratio from perturbation theory of
${\cal B}=\left(\frac{1.42\;\rm GHz}{4\times 14.0\;{\rm(GHz/T)}\times B}\right)^{2}$ (1)
where the numerator is from the 1S hyperfine splitting and the denominator is from the Zeeman splitting when flipping the positron spin. For $B\sim 1$ T, the branching ratio is $\sim 6.4\times 10^{-4}$. Thus, the cooling of the 1Sc population in the ALPHA trap should use less than $\sim 100$ photons to lose less than 10% of the population to spin flip. The $\bar{\rm H}$ can be cooled to steady state with less than $\sim 100$ photons so the spin flip is not a problem as long as the cooling is not overextended. For smaller $B$, the spin flip losses are larger and could become a serious issue below $\sim 0.1$ T.
Figure 1: The splittings of the 1S and 2P states relative to their zero field average. The hyperfine terms of the 2P states are too small to resolve.
@@ III Monte Carlo: no magnetic trap
As discussed in the introduction, laser cooling of trapped $\bar{\rm H}$ is complicated by the changing transition frequency due to the motion through the spatially varying $B$-field. The cooling is also complicated by the fact that the photons only travel in one direction and thus can only cool one component of the velocity. Cooling of all directions relies on the motion through the trap to mix the velocity components. To help with understanding the changes to optical molasses cooling due to the special circumstances of the ALPHA trap, we present in this section results when the restrictions of the trap are removed. Section III.1 reminds the reader of a small energy shift in the laser tuning that arises from the small mass and large transition frequency for $\bar{\rm H}$. Section III.2 describes an algorithm for laser cooling similar in form to that used in the theory of the actual trap but idealizes the physics by having fast mixing of the velocities and neglecting the frequency shift with position; because of the small mass and large transition frequency, this treatment does not lead to the standard optical molasses temperature versus detuning.
@@ III.1 Energy shift
A well known recoil term is usually left out in descriptions of laser cooling, due to its smallness in determining the energy conservation of the transition. For $\bar{\rm H}$ cooling, this term is large enough that it could cause some changes in the results. Consider the energy before and after the photon absorption
@@ III.2 Ideal Optical Molasses
In this section, we describe a method for Monte Carlo simulations of laser cooling of $\bar{\rm H}$ using an ideal optical molassesFoot (2005). We use a method similar to Sec. IV but is idealized by having the velocity mix much faster than the times between photon scattering and by not including the shift in detuning as the $\bar{\rm H}$ moves through the trap. Ideal optical molasses cooling assumes the photon has a wave vector, $\vec{k}$, in a random direction relative to the atom’s velocity, $\vec{v}$. The Doppler effect leads to an effective detuning of
@@
Our Monte Carlo simulation of the ideal optical molasses used the following algorithm: (1) for each of $N$ atoms, randomize the direction of the atom’s velocity, $\vec{v}$, and for that atom compute the effective detuning, Eq. (4); (2) compute the probability that the photon was absorbed using $P=P_{0}/[1+(2\Delta/\Gamma)^{2}]$ with $P_{0}$ a small number of order 0.01; (3) compute a random number; if it is smaller than $P$, then a photon was absorbed; (4) if a photon is absorbed, compute a random emission direction and use Eq. (5) to update that atom’s velocity. If we wanted to simulate the effect of thermalization from elastic collisions between the atoms, we added a step: (5) randomly pick pairs of atoms and rotate their relative velocity to emulate a collision; repeated often enough step (5) leads to a Maxwell-Boltzmann velocity distribution at temperature $k_{B}T=(2/3)\langle KE\rangle$.
@@
Figure 2: For an ideal optical molasses, the steady state temperature using the Monte Carlo method versus the lab frame detuning in units of $\Gamma$. For (a), there is no collision between the atoms to achieve a Maxwell-Boltzmann velocity distribution. The different lines are for different $\bar{\rm H}$ masses to illustrate the role that the small atomic mass plays in the steady state temperature: blue long dash ($1M$), green dash ($2M$), black dot ($4M$), orange dash-dot ($8M$), maroon dash-dot-dot ($16M$), and red solid is the ideal case, Eq. (6). For (b), we compare the effect that elastic collisions leading to thermalization has on the steady state temperature: no collisions [blue long dash ($1M$), green dash ($2M$)], with collisions [orange dash-dot ($1M$), maroon dash-dot-dot ($2M$)].
The results of the simulations are shown in Fig. 2. Figure 2(a) shows the results of the simulations when we did not include collisions between the atoms, skipping step (5) of the algorithm. This figure shows the results when the photon wavelength is kept at 121.6 nm but we artificially change the mass of the $\bar{\rm H}$ from 1 to 16 times the actual mass by factors of 2 in order to illustrate the role that the small atomic mass plays in the change of the steady state temperature from the ideal result of Eq. (6). The blue long-dash curve is for the actual $\bar{\rm H}$ mass. The red solid curve is the ideal case, Eq. (6).
@@ IV Monte Carlo: antihydrogen trap
Some of the features specific for the optical molasses cooling include the path of the laser which is a straight line with a $2.3^{\circ}$ tilt relative to the trap axis taken to be the $z$-direction. Thus, the $\vec{k}$ is a constant. The motion of the $\bar{\rm H}$ through the trap mixes the different velocity components possibly allowing for three dimensional cooling. Since this mixing is relatively slow, the detuning and intensity need to be chosen to limit the heating of the radial motion due the photon emission step. The beam has a waist of 3.48 mm. On the trap axis, $B_{\perp}/B_{z}<0.0003$ while at a radius of 5 mm the $B_{\perp}/B_{z}<0.05$ which means the direction of $\vec{B}$ has little effect. The laser is pulsed with a duration of 10’s of nanoseconds and with repetition rate of 50 Hz. The laser cooling happens over a time scale measured in hours which is possible due to the cryogenic vacuum.
@@
One of the important features is that the detuning of the photon depends on the position in the trap as well as the Doppler shift. The reason for this is the $B$-field changes with position in the trap and the $1S-2P$ resonance frequency increases with increasing $B$-field. For negative detuning, the detuning becomes increasingly negative with increasing $B$-field. Figure 3 shows the magnitude of the $B$-field on the trap axis as a function of the position on the axis relative to the trap center. There are 5 mirror coils and 2 solenoidal coils that are used to shape the $B$-field on the axis. The red solid line is a flat field similar to that used in the 1S2S spectroscopy measurements in Ref. Ahmadi et al (ALPHA collaboration). The blue long dash line gives a more harmonic trap used in some of the simulations below. The green dashed and black dot curves are the slightly dipped and raised traps of Sec. IV.2.
@@
The laser line width, estimated as 60 MHz full-width half-maximum (FWHM), also modestly affects the results since the intrinsic linewidth is approximately 100 MHz. The laser line width is incorporated into the simulation by having each pulse at a slightly different detuning randomly chosen from a Gaussian distribution with a FWHM of 60 MHz. This leads to an effectively broader transition although the effect is not large. We verified that this is a very good approximation (errors less than a couple percent) by comparing to the results from numerically solving the Optical Bloch equations for the same parameters. In all of the figures and results, the linewidth, $\Gamma$, is the natural radiative linewidth of the 2P state.
@@ IV.3 Cooling in a nearly harmonic trap
The harmonic trap, blue long dash curve in Fig. 3, has only a tiny region where the detuning has smallest magnitude. This suggests that the problems with the photon scattering rate being higher than the coordinate mixing rate will be less. The actual case is that the harmonic trap has much smaller mixing rate between the coordinates. We find that the temperature differences are much more extreme than the previous section. For example, at $\Delta_{0}=-3\Gamma$ and 2 nJ, $T_{z}\simeq 6$ mK and $T_{r}\simeq 24$ mK for the harmonic trap while the flat trap has $\simeq 6$ and $\simeq 9$ mK. Thus, nearly harmonic traps appear to be a poor choice for laser cooling of only the axial motion. Similarly, making a much larger hump in the previous section also leads to less effective cooling because the two side wells become more harmonic like.
@@ V Simultaneous cooling 1Sc and 1Sd
Unfortunately, problems arise when laser cooling both populations and trying to achieve the same distributions for each. The 1Sc-2Pa transition frequency is 675 MHz larger at $B\sim 1$ T than the 1Sd-2Pa. Detuning to -250 MHz for the 1Sd-2Pa transition gives -925 MHz detuning for the 1Sc-2Pa transition. Thus, the cooling is extremely slow for the 1Sc states and the final distribution is at a substantially larger temperature with, essentially, uncooled $\bar{\rm H}$s.Baker et al (ALPHA collaboration) In principle, the 1Sc and 1Sd states could be mixed using microwaves with frequencies less than 1 GHz, but such long wavelength light does not propagate down the $\sim 22$ mm radius ALPHA trap; this would require a special resonator and microwave input in the next generation trap.
@@ VI Summary
Although we have not simulated the following modifications, the outcomes seem to follow from the results presented above. One possible modification would be to increase the angle of the laser with respect to the trap axis. While this is not possible in the current ALPHA experiment, a larger angle (say between 20-40) would directly cool two of the spatial coordinates (for example, $x$ and $z$) and only require mixing of the coordinate perpendicular to the plane defined by the laser and the trap axis. Another possibility is to use stronger laser pulses with larger detuning early in the cooling cycle and then transition to weaker pulses and smaller detuning near the end of cooling; this should lead to faster cooling at early times and lower temperatures at later times.